INFRARED EMISSION SPECTRUM OF He_{2} OBSERVED BY A PULSED DISCHARGE

YUKIO HOSAKI, and \underline{KENTAROU KAWAGUCHI} , Faculty of Science, Okayama University, Okayama 700-8530, Japan; SVATOPLUK CIVIS, J. Heyrovsky Institute of Physical Chemistry, Academy of Science of the Czech Republic, Dolejskova 3, 18223 Prague 8, Czech Republic.

The He_{2} molecule is known to have many electronic transitions between Rydberg states in visible and ultraviolet regions. However, in infrared region only two bands b^{3}Π_{g} - a^{3}Σ_{u}^{+} and B^{1}Π_{g} - A^{1}Σ_{u}^{+} have been studied by using a DC discharge method so far. Recently we developed a time resolved Fourier transform spectroscopy with high resolution Bruker IFS 120 HR by using micro controller SX^{a}. In the present study, the FT system was applied to infrared emission spectroscopy of He_{2} which was produced by a pulsed discharge in He with pulse width of 20 μsec and 1 A peak-to-peak current. In the 1800-10000 cm^{−1} region, many electronic transitions have been observed in addition to the previously reported two bands. From observed time profiles of emission spectra, Rydberg states with higher energy than the b state are produced efficiently in afterglow plasma after termination of the discharge. A least-squares analysis was carried out for the h^{3}Σ_{u}^{+} - g^{3}Σ_{g}^{+} and g^{3}Σ_{g}^{+} - d^{3}Σ_{u}^{+} bands in the 3200 cm^{−1} region to determine the molecular constants. A transition from an un-identified state to the d^{3}Σ_{u}^{+} state has been observed with irregular Pand R- branch intensities.

aK. Kawaguchi, O. Baskakov, Y. Hosaki, Y. Hama, and C. Kugimiya, Chem. Phys. Lett. \underline{369}, 293(2003)